. You want to buy a car, and a local bank will lend you $15,000. The loan will be fully amortized over 3 years and the nominal interest rate will be 9%. Construct an amortization table showing the fixed payment, the beginning year principal payment, the interest payment, and the ending principal balance for each year. Note that you will get a partial credit if you just show numbers
Answer #1

Yearly Loan Payment = Amount / Discount Factor or P = A / D

Number of Periodic Payments (n) = Payments per year times number of years= 3

Periodic Interest Rate (i) = Annual rate divided by number of payment periods= 9%

Discount Factor (D) = {[(1 + i) ^n] - 1} / [i(1 + i)^n]

Here, Discount factor= (((1+0.09)^3 -1)/(0.09(1+0.09)^3)

D= 2.53

Amount= 15000

Loan payment = 15000/2.53

= 5925.821

So the amortization table is as follow

Year

Opening balance

Interest

Interest

Principal= Annual payment - interest

Principal Paid

Total amount paid

Balance= Opening balance- principal paid

Closing Balance

1

15,000.00

15000*9%=

1,350.00

5925.82-1350=

4,575.82

5,925.82

10,424.18

10,424.18

2

10,424.18

10424.18*9%

938.18

5925.82-938.17=

4,987.65

5,925.82

7,022.26

5,436.53

3

5,436.53

5436.53*9%

489.29

5925.82-1350=

5,436.53

5,925.82

2,624.27

0.00